OpenStudy (anonymous):

The company offers you a 42-day term loan with an interest rate of 6.8% . What is the effective interest rate on this loan? [express your answer in percentage rounded to two decimal places]

OpenStudy (rane):

r = [1 + (i/n)]n - 1 = (1 + (6.8/42) 42 - 1 ) = 7.031%